87.48 g of an unknown compound composed of carbon, hydrogen and oxygen contains 41.17 g C and 5.18 g H. Give the empirical formula of this compound.

C4H5O2

C3H5O2

C3H5O4

C3H5O3

C4H6O3
Answer #1

Let's get the composition of Oxygen:

mass O = 87.48 - 41.17 - 5.18 = 41.13 g

Now, let's get the moles:

C = 41.17 / 12 = 3.43

H = 5.18 / 1 = 5.18

O = 41.13 / 16 = 2.57

Now, the relation between then to get the number of atoms:

C = 3.43 / 2.57 = 1.33

H = 5.18 / 2.57 = 2.01

O = 2.57 / 2.57 = 1

The empirical formula should be:

C1.33H2O

However this does not match the options...at first.

However, if we just have the molecular weight of this compound, we can calculate the molecular formula (which I think is the real question here).

the composition of C, H and O in this is:

%C = 41.17/87.48 * 100 = 47.1%

%H = 5.18/87.48 * 100 = 5.9%

%O = 41.13/87.48 * 100 = 47%

So, if we calculate the percent for every option, the only one that match would be the last compound, which mean that the formula is C4H6O3 but this would be molecular formula. The empirical is the above.

Hope this helps